OpenStudy (anonymous):

Solve exactly in terms of ln or log. be sure to check solutions: 2^2x + 2^x - 12 = 0

OpenStudy (anonymous):

solve as a quadratic in \[2^x\] say put \[z=2^x\] and then solve \[z^2+2z-12=0\] \[z=-1+\sqrt{13}\] \[2^x=-1+\sqrt{13}\] \[x=\log_2(-1+\sqrt{13})\]